OpenStudy (anonymous):

Solve: -1/2x + 2 = −x + 7

OpenStudy (saifoo.khan):

@GOODMAN

OpenStudy (anonymous):

Subract 7 from both sides. -1/2x-5=-x Add 1/2x -5=-1/2x Solve from there.